A long, thin rod (length = 3.0 m) lies along the x axis, with its midpoint at the origin. In a vacuum, a +6.0C point charge is fixed to one end of the rod, and a -6.0C point charge is fixed to the other end. Everywhere in the x, y plane there is a constant external electric field (magnitude 3.0 × 103 N/C) that is perpendicular to the rod. With respect to the z axis, find the magnitude of the net torque applied to the rod.
Answer #1

answer) we know the formula

=F*l

F=qE

the force create a counter clockwise torque , so they add up and the torque will be produced through the center of the rod because z axis is the middle point of the the rod of the length.L

so l=L/2

so =qE*L/2+qE*L/2=qEL=6C*3*103N/C*3m=54*103N.m

so the answer is 54000 N.m or 54*103N.m or 5.4*104 N.m